# Sentiment-Aware Stock Price Prediction with Transformer and LLM-Generated Formulaic Alpha

### v1

@@ 4.3 Effect of Sentiment Integration on Prediction Accuracy
Table 6 compares the performance of the Transformer model with and without sentiment information. The experiments are run ten times for each stock. Across the five companies, the incorporation of sentiment features leads to noticeable improvements in prediction accuracy, as reflected in consistently lower MSE. Additionally, standard deviations are generally reduced in the sentiment-enhanced setting, suggesting more stable and robust model behavior. This indicates that sentiment information not only contributes to better predictive performance, but also helps mitigate variability across different runs. Although the magnitude of improvement varies by company, the consistent trend reinforces the value of leveraging sentiment from both target firms and their related entities in stock price forecasting.
@@
Table 6: Transformer Performance with and without Sentiment Information (Mean and Standard Deviation of MSE over 10 runs)
@@ 5.2 Further Evaluation of LLM-Generated Alphas
Alpha 1 -0.0321 0.0345 -0.0419 -0.0105 -0.0267
Alpha 2 -0.0114 0.0307 -0.0226 0.0193 0.0169
Alpha 3 -0.0152 0.0330 0.0009 0.0322 0.0092
Alpha 4 0.0210 0.0124 -0.0018 -0.0128 0.0142
Alpha 5 -0.0028 0.0191 0.0551 -0.0149 0.0140
@@ 6 Conclusion
This study proposes a novel framework that integrates LLM-generated formulaic features with a Transformer model to forecast stock movement. The formulaic alphas are derived from basic inputs, such as stock features, technical indicators, and sentiment scores. For sentiment data, both target companies and their related companies are considered. Related companies are initially extracted using NER, based on their frequency of occurrence in news articles about the target companies. The most relevant ones, those strongly correlated with stock movement, are further selected by the LLM and incorporated into the formulaic alphas. For comparison of performance, models including LSTM, TCN, SVR, and Random Forest are used. The results highlight the predictive strength of the LLM-generated features, with the Transformer model achieving the best performance in most scenarios.
For future study, it can be explored how the number of LLM-generated formulaic alphas will affect prediction performance. Different prompt designs to generate formulaic alphas and different LLMs can be used for comparison.

### v2

@@ 4.3 Effect of Sentiment Integration on Prediction Accuracy
Table 9 compares the performance of the Transformer model with and without sentiment information. The experiments are run ten times for each stock to consider the feature randomness because each run generates a new alpha set. Across the five companies, the incorporation of sentiment features leads to noticeable improvements in prediction accuracy, as reflected in consistently lower MSE. Additionally, standard deviations are generally reduced in the sentiment-enhanced setting, suggesting more stable and robust model behavior. This indicates that sentiment information not only contributes to better predictive performance, but also helps mitigate variability across different runs. Although the magnitude of improvement varies by company, the consistent trend reinforces the value of leveraging sentiment from both target firms and their related entities in stock price forecasting.
@@
Table 9: Transformer performance with and without sentiment information across 10 different alpha sets, reported as mean and standard deviation of MSE
@@
Table 10: MSE comparison of stock price prediction models using a fixed alpha set over 10 runs (mean and standard deviation), with the best (lowest MSE) in each column in bold and the second best underlined.
Model Apple HSBC Pepsi Tencent Toyota
Transformer 0.00056 ($3.04\times 10^{-4}$) 0.00040 ($1.72\times 10^{-4}$) 0.00020 ($1.32\times 10^{-4}$) 0.00047 ($4.06\times 10^{-4}$) 0.00049 ($1.84\times 10^{-4}$)
Informer 0.00054 ($2.82\times 10^{-4}$) 0.00074 ($2.54\times 10^{-4}$) 0.00046 ($2.37\times 10^{-4}$) 0.00094 ($3.35\times 10^{-4}$) 0.00052 ($2.24\times 10^{-4}$)
LSTM 0.00298 ($1.06\times 10^{-4}$) 0.00071 ($1.48\times 10^{-4}$) 0.00170 ($2.47\times 10^{-4}$) 0.00172 ($1.22\times 10^{-4}$) 0.00128 ($3.98\times 10^{-4}$)
TCN 0.00178 ($1.03\times 10^{-4}$) 0.00054 ($0.35\times 10^{-4}$) 0.00266 ($5.39\times 10^{-4}$) 0.00186 ($3.99\times 10^{-4}$) 0.00302 ($6.74\times 10^{-4}$)
SVR 0.00479 0.00157 0.00436 0.00251 0.00481
Random Forest 0.00186 ($0.09\times 10^{-4}$) 0.00044 ($0.02\times 10^{-4}$) 0.00140 ($0.15\times 10^{-4}$) 0.00103 ($0.05\times 10^{-4}$) 0.00050 ($0.05\times 10^{-4}$)
XGBoost 0.00189 ($0.72\times 10^{-4}$) 0.00049 ($0.21\times 10^{-4}$) 0.00151 ($0.49\times 10^{-4}$) 0.00116 ($0.29\times 10^{-4}$) 0.00051 ($0.15\times 10^{-4}$)
Table 11: Variance decomposition of MSE across repeated experiments based on results of Table 10. The total variance is decomposed into between-model variance and within-model variance based on 10 independent runs for each stock. The percentage columns indicate the relative contribution of each component to the total variance.
Stock Between-Model Variance Within-Model Variance Total Variance Between (%) Within (%) Apple $2.17\times 10^{-6}$ $3.32\times 10^{-8}$ $2.20\times 10^{-6}$ 98.49 1.51 HSBC $1.64\times 10^{-7}$ $1.96\times 10^{-8}$ $1.84\times 10^{-7}$ 89.34 10.66 Pepsi $1.98\times 10^{-6}$ $7.13\times 10^{-8}$ $2.05\times 10^{-6}$ 96.53 3.47 Tencent $4.69\times 10^{-7}$ $7.53\times 10^{-8}$ $5.45\times 10^{-7}$ 86.17 13.83 Toyota $2.87\times 10^{-6}$ $1.16\times 10^{-7}$ $2.99\times 10^{-6}$ 96.11 3.89
4.4 Training Stability of Models with a Fixed Alpha Set
Table 10 illustrates the performance comparison of different models using a fixed alpha set over ten runs, reported as mean and standard deviation. In this study, SVR adopts a linear kernel setup, so the results of SVR are deterministic.
The results show clear differences across models when a fixed alpha set is used. Overall, attention-based models deliver more accurate predictions than the other approaches. The Transformer model performs consistently well across most stocks, suggesting that its attention structure is effective in capturing time dependencies under this setting. Informer also shows competitive performance, especially for companies such as Apple, but its results are relatively less stable across different stocks.
In contrast, recurrent and convolution-based models, such as LSTM and TCN, tend to produce higher prediction errors. This indicates that these models may have difficulty modeling long-range patterns between different alphas. Traditional machine learning methods show mixed behavior. Tree-based models such as Random Forest and XGBoost perform reasonably well for some stocks but lack consistency across all cases, while SVR generally yields weaker results.
These findings suggest that attention-based architectures are more suitable for stock price forecasting when the same alpha set is applied across models. Their ability to focus on relevant time steps likely helps improve both accuracy and stability, whereas other models appear more sensitive to feature limitations.
Besides, to further examine the robustness of the empirical results, we conduct a variance decomposition analysis based on repeated experiments. Specifically, the total variance of the MSE is decomposed into two components: (i) between-model variance, which captures performance differences attributable to model choice, and (ii) within-model variance, which reflects variability arising from training randomness such as parameter initialization and stochastic optimization. This analysis allows us to assess whether the observed performance differences are primarily driven by model selection or by random fluctuations across runs.
Table 11 presents the variance decomposition results for all five stocks. Across all cases, the between-model variance accounts for the dominant share of the total variance, ranging from 86.17% to 98.49%. In contrast, the within-model variance remains relatively small, indicating that the forecasting performance of each model is stable across repeated runs. These findings suggest that the differences observed in the MSE comparison are mainly driven by inherent model characteristics rather than training randomness.
@@ 5.2 Further Evaluation of LLM-Generated Alphas
Alpha 1 0.0321-0.0321 0.0345 0.0419-0.0419 0.0105-0.0105 0.0267-0.0267
Alpha 2 0.0114-0.0114 0.0307 0.0226-0.0226 0.0193 0.0169
Alpha 3 0.0152-0.0152 0.0330 0.0009 0.0322 0.0092
Alpha 4 0.0210 0.0124 0.0018-0.0018 0.0128-0.0128 0.0142
Alpha 5 0.0028-0.0028 0.0191 0.0551 0.0149-0.0149 0.0140
@@ 6 Conclusion
This study proposes a novel framework that integrates LLM-generated formulaic features with a Transformer model to forecast stock movement. The formulaic alphas are derived from basic inputs, such as stock features, technical indicators, and sentiment scores. For sentiment data, both target companies and their related companies are considered. Related companies are initially extracted using NER, based on their frequency of occurrence in news articles about the target companies. The most relevant ones, those strongly correlated with stock movement, are further selected by the LLM and incorporated into the formulaic alphas. For comparison of performance, models including Informer, LSTM, TCN, SVR, Random Forest, and XGBoost are used. The results highlight the predictive strength of the LLM-generated features.
This study has several limitations. First, the analysis is conducted on only five companies and within a single evaluation period. Future studies may consider a larger set of stocks and multiple time periods to improve generalizability. In addition, more recent and advanced time-series architectures such as N-BEATS can be included as baseline models for comparison. Besides, it can be explored how the number of LLM-generated formulaic alphas will affect prediction performance. Different prompt designs to generate formulaic alphas and different LLMs can be used for comparison.